OpenStudy (anonymous):

another help!!! how do i find the x and y intercept from a graph i.e. y=cos(5x/6+pi/3)

OpenStudy (anonymous):

**3cos

OpenStudy (bahrom7893):

x - intercept means y=0 y - intercept means x=0

OpenStudy (bahrom7893):

cos(5x/6+pi/3)=0 so solve for x

OpenStudy (bahrom7893):

y=cos(pi/3) is the y intercept

OpenStudy (bahrom7893):

y=1/2 is the y intercept

OpenStudy (bahrom7893):

wait sorry u said 3Cos

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

y=3cos(5x/6+pi/3)

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

Hmm x intercepts are tough sort of.. http://www.wolframalpha.com/input/?i=Solve%283*cos%285*x%2F6%2Bpi%2F3%29%3D0%2Cx%29

OpenStudy (bahrom7893):

Y intercepts: y=3cos(pi/3) = 1.5

OpenStudy (bahrom7893):

* there's only one y intercept btw

OpenStudy (bahrom7893):

Please click on Good Answer if I helped, I really want to get to the next level, thanks =)!

OpenStudy (anonymous):

thanx!!!

OpenStudy (bahrom7893):

np

OpenStudy (anonymous):

its wrong.. cos(5x/6+pi/3) is not equal to cos(5X/6)+cos(pi/3)

OpenStudy (anonymous):

3cos***